OpenStudy (anonymous):

Simplify.7 square root of 3 end root minus 4 square root of 6 end root plus square root of 48 end root minus square root of 54 11 square root of 6 end root minus 7 square root of 12 11 square root of 3 end root minus 7 square root of 6 negative 3 square root of 9 4 square root of 9

OpenStudy (anonymous):

@lizz123

OpenStudy (anonymous):

You can only add roots that have the same number inside. So start by simplifying the bigger roots (sqrt48 and sqrt54).

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What did you get when you simplified?

OpenStudy (anonymous):

idk @jwhite12 i believe the answer is c o. d

OpenStudy (anonymous):

\[\sqrt{48}=\sqrt{16}\sqrt{3}=4\sqrt{3}\] \[\sqrt{54}=\sqrt{9}\sqrt{6}=3\sqrt{6}\] So then the problem is: \[7\sqrt{3} - 4\sqrt{6} + 4\sqrt{3} - 3\sqrt{6}\] Add or subtract the outside numbers of the like radicals.

OpenStudy (anonymous):

@SolomonZelman @sleepyjess

OpenStudy (sleepyjess):

Just do what @jwhite12 is telling you.

OpenStudy (anonymous):

idk how to do that @sleepyjess

OpenStudy (sleepyjess):

\(7\sqrt3 + 4\sqrt3\)

OpenStudy (sleepyjess):

what is 7+4

OpenStudy (anonymous):

11 @sleepyjess

OpenStudy (anonymous):

the answer is b @sleepyjess

OpenStudy (sleepyjess):

Ok, but do you understand?

OpenStudy (anonymous):

yes @sleepyjess

OpenStudy (anonymous):

Which statement is true about the product square root of 2(3square root of 2 + square root of 14)? It is rational and equal to 12. It is rational and equal to 18. It is irrational and equal to 6 + 2square root of 18. It is irrational and equal to 3 +square root of 18.

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

I think that @jwhite12 is better at this than I am

OpenStudy (anonymous):

i think he is asleep or not paying attention to us tbh

OpenStudy (anonymous):

Can you type the question in the equation editor?

OpenStudy (sleepyjess):

^^^^^^ that would also make it easier to help

OpenStudy (anonymous):

um sure

OpenStudy (anonymous):

\[\sqrt{2}(\sqrt[3]{2} +\sqrt{18}\]

OpenStudy (anonymous):

is that better??? @sleepyjess @jwhite12

OpenStudy (anonymous):

Is it \[3\sqrt{2}\] or \[\sqrt[3]{2}\]

OpenStudy (anonymous):

Start by simplifying the \[\sqrt{18}\] Do you know how to do that?

OpenStudy (anonymous):

2nd one and no

OpenStudy (anonymous):

@jwhite12 @sleepyjess

OpenStudy (anonymous):

You have to break 18 into factors where one is a perfect square. So 18 breaks into 9 and 2. Then you square root the 9 to get 3, and that's your new outside number, keeping the 2 inside. So...

OpenStudy (anonymous):

\[\sqrt{18}=\sqrt{9}\sqrt{2}=3\sqrt{2}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so d @jwhite12

OpenStudy (anonymous):

Are you sure it's not written as \[3\sqrt{2}\] in the original problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think that 3 is supposed to be in front of the radical, not the index. I know it looks a little higher, but it may be the equation editor your teacher is using. If it were the index if would be inside the radical (check mark thing). And that's not the type of question you've been having anyway. AND if the index is 3, then none of those answers are right. So I'm going to assume it's outside...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

That gives us: \[\sqrt{2}(3\sqrt{2}+3\sqrt{2})\] We have "like radicals" inside the parentheses, so we can add the outside numbers...

OpenStudy (anonymous):

Then we have \[\sqrt{2}(6\sqrt{2})\]

OpenStudy (anonymous):

We can multiply the numbers inside the radical and keep the 6 outside, so we get \[6\sqrt{4}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you know what that simplifies to?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Do you know the square root of 4?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Good, so that problem becomes 6 times 2.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

That's it!

OpenStudy (anonymous):

thank you d oyou mind checking if my answer are correct for my other questions

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Simplify. square root of 5 open parentheses 10 minus 4 square root of 2 close parentheses 14 15 square root of 2 5 square root of 2 end root minus 4 square root of 10 None of the above

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

@jwhite12

OpenStudy (anonymous):

Is it \[\sqrt{5}(10-4\sqrt{2})\] ?

OpenStudy (anonymous):

yes @jwhite12

OpenStudy (anonymous):

please just tell me is it correct

OpenStudy (anonymous):

yes it's d